# Reinforcement Learning for Reasoning in Small LLMs: What Works and What Doesn’t

### v1

@@ Abstract
Enhancing the reasoning capabilities of large language models (LLMs) typically relies on massive computational resources and extensive datasets, limiting accessibility for resource-constrained settings. Our study investigates the potential of reinforcement learning (RL) to improve reasoning in small LLMs, focusing on a 1.5-billion-parameter model, DeepSeek-R1-Distill-Qwen-1.5B, under strict constraints: training on 4 NVIDIA A40 GPUs (48 GB VRAM each) within 24 hours. Adapting the Group Relative Policy Optimization (GRPO) algorithm and curating a compact, high-quality mathematical reasoning dataset, we conducted three experiments to explore model behavior and performance. Our results demonstrate rapid reasoning gains—e.g., AMC23 accuracy rising from 63% to 80% and AIME24 reaching 46.7%, surpassing o1-preview—using only 7,000 samples and a \$42 training cost, compared to thousands of dollars for baseline models. However, challenges such as optimization instability and length constraints emerged with prolonged training. These findings highlight the efficacy of RL-based fine-tuning for small LLMs, offering a cost-effective alternative to large-scale approaches. We release our code and datasets as open-source resources, providing insights into trade-offs and laying a foundation for scalable, reasoning-capable LLMs in resource-limited environments. All are available at https://github.com/knoveleng/open-rs.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have significantly advanced the pursuit of artificial general intelligence (AGI), with models such as GPT-4o (OpenAI, 2024a), Claude 3.5 Sonnet (Anthropic, 2024), and Gemini 1.5 (Google, 2024) demonstrating unprecedented capabilities. A pivotal aspect of this progress is the integration of post-training techniques into the training pipeline. These methods—including supervised fine-tuning (SFT) and reinforcement learning (RL)—enhance reasoning accuracy, align models with societal values, and adapt them to user preferences, all while demanding fewer computational resources than pre-training (OpenAI, 2024b). A notable innovation in this domain is OpenAI’s o1 series, which leverages inference-time scaling through extended Chain-of-Thought (CoT) reasoning to achieve remarkable performance in mathematics, coding, and scientific reasoning tasks (OpenAI, 2024b). However, despite these breakthroughs, scaling reasoning capabilities at test time remains a persistent challenge for the broader research community, largely due to limited access to proprietary methodologies and resources.
Efforts to bolster LLM reasoning have explored diverse strategies. Process-based reward models (Uesato et al., 2022; Lightman et al., 2023a; Wang et al., 2023) guide models toward structured problem-solving, while RL approaches (Kumar et al., 2024) optimize performance through feedback-driven learning. Search algorithms, such as Monte Carlo Tree Search (MCTS) and Beam Search, have also been employed to enhance reasoning depth (Feng et al., 2024; Xin et al., 2024; Trinh et al., 2024). Although these methods have driven incremental gains, they fall short of the general reasoning prowess exhibited by the o1 series. Recently, the DeepSeek-R1 model (DeepSeek-AI, 2025) has emerged as a competitive alternative, utilizing RL with the Group Relative Policy Optimization (GRPO) algorithm. Built on the 671-billion-parameter DeepSeek-V3, DeepSeek-R1 matches o1’s reasoning performance (DeepSeek-AI, 2025). Yet, the sheer scale and computational demands of such models—often exceeding hundreds of billions of parameters—render them impractical for self-hosting by most organizations outside major technology firms, limiting their broader adoption.
In contrast, small LLMs, typically ranging from 1 to 10 billion parameters, present a resource-efficient alternative with potential for widespread deployment. Previous studies have demonstrated the feasibility of enhancing small LLMs through RL-based fine-tuning inspired by DeepSeek-R1 (Luo et al., 2025; Team, 2025b). However, these efforts often rely on expansive datasets (hundreds of thousands to millions of samples) or incur significant computational costs, undermining their accessibility for resource-constrained settings. This tension motivates two central research questions:
@@
Our findings illuminate the promise of RL-based methods to enhance small LLMs’ reasoning capabilities, achieving competitive performance with minimal resources (Figure 1). Simultaneously, they reveal critical challenges—such as data efficiency, optimization stability, and length constraints—that must be addressed to fully realize this potential. These insights lay the groundwork for developing lightweight, reasoning-capable LLMs suitable for resource-constrained environments, advancing the democratization of advanced AI technologies.
@@ 2 Methodology
2.1 High-Quality Dataset Curation
To minimize training costs while maximizing reasoning performance, we curate a compact, high-quality dataset tailored to mathematical reasoning. This dataset is derived from two existing sources: the s1 dataset (Muennighoff et al., 2025) and the DeepScaleR dataset (DeepSeek-AI, 2025). By filtering and refining these datasets, we ensure that our training data is both relevant and challenging, enabling efficient learning for small LLMs.
@@ 2.1 High-Quality Dataset Curation
The s1 dataset (Muennighoff et al., 2025) is a general-purpose reasoning corpus comprising 59,029 questions sourced from diverse domains, including NuminaMATH (LI et al., 2024), AIME problems (1983–2021), OlympicArena (Huang et al., 2024), OmniMath (Gao et al., 2024), AGIEval (Zhong et al., 2023), probability questions from Stanford University’s Statistics Department PhD Qualifying Exams (https://statistics.stanford.edu), and brain-teasers from PuzzledQuant (https://www.puzzledquant.com). Although the dataset spans multiple disciplines—such as Astronomy, Biology, Chemistry, Computer Science, Geography, Mathematics, and Physics—our focus is exclusively on mathematical reasoning.
@@
The DeepScaleR dataset (Luo et al., 2025) contains 40,315 mathematics-specific questions drawn from AIME (1984–2023), AMC (prior to 2023), Omni-MATH, and the Still dataset. Unlike the s1 dataset, DeepScaleR is pre-filtered to focus solely on mathematics, with redundant questions removed and solutions extracted from raw text using retrieval-augmented generation (RAG) and advanced LLMs like Gemini-1.5-Pro-002. To further refine this dataset, we apply Qwen2.5-Math-7B-Instruct to exclude easy questions, reducing the set to 21,044 examples – open-deepscaler dataset. We opt for Qwen2.5-Math-7B-Instruct over DeepSeek-R1-Distill-Qwen-1.5B—used for the s1 dataset—to introduce diversity in filtering criteria and avoid excessive overlap between the two datasets.
@@
2.2 Reinforcement Learning Algorithm
To train small LLMs efficiently, we adopt the Group Relative Policy Optimization (GRPO) algorithm Shao et al. (2024), as utilized in DeepSeek-AI (2025). GRPO eliminates the need for a separate critic model—typically as large as the policy model—by estimating baselines from group scores, thereby reducing computational overhead. For each question $q$, GRPO samples a group of $G$ outputs $\{o_{1},o_{2},\dots,o_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$ and optimizes the policy $\pi_{\theta}$ by maximizing the following objective:
$\begin{split}\mathcal{J}_{\text{GRPO}}(\theta)&=\mathbb{E}_{[q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)]}\\ &\frac{1}{G}\sum_{i=1}^{G}\left(\min\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}A_{i},\text{clip}\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)},1-\epsilon,1+\epsilon\right)A_{i}\right)-\beta\mathbb{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\right),\end{split}$ (1)
@@ 3 Experiments
To address the research questions outlined in Section 1—namely, how reinforcement learning (RL) can enhance the reasoning abilities of small large language models (LLMs) and what practical insights emerge under computational constraints—we design three experiments to analyze the training behavior of small LLMs. These experiments aim to provide empirical evidence of performance improvements and offer actionable guidance for future research and industrial applications.
3.1 Experimental Setup
We select DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI, 2025) as our base model for training. This 1.5-billion-parameter model, distilled from larger architectures, is chosen for its balance of efficiency and reasoning potential. Notably, we bypass the supervised fine-tuning (SFT) phase—typically a precursor to RL for performance enhancement (Chu et al., 2025)—hypothesizing that the model’s pretraining is sufficient to leverage RL directly. For the RL phase, we employ the Group Relative Policy Optimization (GRPO) algorithm, as detailed in Section 2.2, due to its computational efficiency.
Training is conducted on a cluster of 4 NVIDIA A40 GPUs (48GB VRAM each), imposing constraints that limit us to sampling 6 outputs per step with a maximum completion length of 4096 tokens. To facilitate this, we adapt open-r1 (Face, 2025), an open-source reproduction of DeepSeek-R1 by the Hugging Face team, customizing it to align with our objectives. The training phase is restricted to 1 epoch, completed within a 24-hour window, reflecting real-world resource limitations. Hyperparameters and additional configurations are detailed in Appendix E.
3.2 Benchmark Datasets
To evaluate the reasoning capabilities of our small LLM, we choose five mathematics-focused benchmark datasets: AIME24  [^1] https://huggingface.co/datasets/AI-MO/aimo-validation-aime, MATH-500 (Lightman et al., 2023b; Hendrycks et al., 2021), AMC23  [^2] https://huggingface.co/datasets/AI-MO/aimo-validation-amc, Minerva (Lewkowycz et al., 2022b) and OlympiadBench (He et al., 2024). Details of the datasets are provided in Appendix C.
3.3 Baseline Models
To contextualize our results, we compare our trained model against a range of baselines: Llama-3.1-70B-Instruct (AI, 2024a), o1-preview (AI, 2024b), Qwen-2.5-Math-7B-Instruct (Yang et al., 2024), rStar-Math-7B (Guan et al., 2025), Eurus-2-7B-PRIME, Qwen2.5-7B-SimpleRL (Zeng et al., 2025) (Cui et al., 2025), DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI, 2025), DeepScaleR-1.5B-Preview (Luo et al., 2025), Still-3-1.5B-Preview (Team, 2025b).
This selection enables a robust comparison across model sizes, training methodologies, and reasoning strategies, highlighting the efficacy of our approach for small LLMs. Details of the baselines are provided in Appendix D.
3.4 Evaluation Metric
@@ 3.5 Process and Results
3.5.1 Experiment 1: Impact of High-Quality Data
In Experiment 1, we train the DeepSeek-R1-Distill-Qwen-1.5B model using the open-s1 dataset (18,615 samples) from Section 2.1, with a maximum completion length of 4096 tokens. We employ accuracy and format rewards, as described in Section 2.2. Although the full dataset corresponds to approximately 1500 global steps for one epoch, computational constraints (24-hour limit on 4x A40 GPUs) restrict training to 500 global steps.
@@ 3.5.1 Experiment 1: Impact of High-Quality Data
3.5.2 Experiment 2: Balancing Easy and Hard Problems
Building on Experiment 1, we hypothesize that mixing easier problems with challenging ones could stabilize training and reduce completion lengths. We construct a dataset of 7000 samples: 3000 from open-s1, 3000 from open-deepscaler, and 1000 easier problems from the raw DeepScaleR dataset (Section 2.1). The maximum completion length is reduced to 3584 tokens, retaining accuracy and format rewards.
@@ 3.5.2 Experiment 2: Balancing Easy and Hard Problems
3.5.3 Experiment 3: Controlling Length with Cosine Reward
Experiment 3 uses the same 7000-sample dataset as Experiment 2, but replaces the accuracy reward with a cosine reward to better control output length, as outlined in Section 2.2. We also append an instruction to the system prompt: “Reply in English only, do not use other languages”, avoiding a computationally expensive language reward function. The maximum completion length remains 3584 tokens.
@@ 3.5.3 Experiment 3: Controlling Length with Cosine Reward
3.5.4 Overall Comparison
We select checkpoints at 100, 50, and 50 global steps from Experiments 1, 2, and 3, naming them Open-RS1, Open-RS2, and Open-RS3 (R for Reasoning, S for Small), respectively. These are evaluated against baselines from Section 3.3 across benchmarks from Section 3.2, using zero-shot pass@1 (Table 1).
@@ 1.5B Models
Table 1: Zero-shot pass@1 performance across benchmarks. Bold indicates the highest score per benchmark. Dashes (–) denote unavailable official scores. Scores for o1-preview are sourced from AI (2024b); others from Zeng et al. (2025); Luo et al. (2025). Our models are evaluated using the lighteval package Fourrier et al. (2023).
@@
We further compare training costs [^3] Cost estimates are calculated based on pricing from https://www.runpod.io/pricing. and data efficiency (Tables 2 and 3, and Figure 1). Our approach, using 7000 samples with 6 outputs per step (42,000 total samples), costs approximately \$42 on 4x A40 GPUs over 24 hours. In contrast, 7B models like Qwen2.5-7B-SimpleRL (\$1633) and Eurus-2-7B-PRIME (\$1088) and 1.5B models like DeepScaleR-1.5B-Preview (\$3629) and Still-3-1.5B-Preview (\$2268) require significantly more resources and data (e.g., 40k × 16 samples for DeepScaleR).

### v2

@@ Abstract
Enhancing the reasoning capabilities of large language models (LLMs) typically relies on massive computational resources and extensive datasets, limiting accessibility for resource-constrained settings. Our study investigates the potential of reinforcement learning (RL) to improve reasoning in small LLMs, focusing on a 1.5-billion-parameter model, DeepSeek-R1-Distill-Qwen-1.5B, under strict constraints: training on 4 NVIDIA A40 GPUs (48 GB VRAM each) within 24 hours. Adapting the Group Relative Policy Optimization (GRPO) algorithm and curating a compact, high-quality mathematical reasoning dataset, we conducted three experiments to explore model behavior and performance. Our results demonstrate rapid reasoning gains - e.g., AMC23 accuracy rising from 63% to 80% and AIME24 reaching 46.7%, surpassing o1-preview - using only 7,000 samples and a \$42 training cost, compared to thousands of dollars for baseline models. However, challenges such as optimization instability and length constraints emerged with prolonged training. These findings highlight the efficacy of RL-based fine-tuning for small LLMs, offering a cost-effective alternative to large-scale approaches. We release our code and datasets as open-source resources, providing insights into trade-offs and laying a foundation for scalable, reasoning-capable LLMs in resource-limited environments. All are available at https://github.com/knoveleng/open-rs.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have significantly advanced the pursuit of artificial general intelligence (AGI), with models such as GPT-4o (OpenAI 2024a), Claude 3.5 Sonnet (Anthropic 2024), and Gemini 1.5 (Google 2024) demonstrating unprecedented capabilities. A pivotal aspect of this progress is the integration of post-training techniques into the training pipeline. These methods - including supervised fine-tuning (SFT) and reinforcement learning (RL) - enhance reasoning accuracy, align models with societal values, and adapt them to user preferences, all while demanding fewer computational resources than pre-training (OpenAI 2024b). A notable innovation in this domain is OpenAI’s o1 series, which leverages inference-time scaling through extended Chain-of-Thought (CoT) reasoning to achieve remarkable performance in mathematics, coding, and scientific reasoning tasks (OpenAI 2024b). However, despite these breakthroughs, scaling reasoning capabilities at test time remains a persistent challenge for the broader research community, largely due to limited access to proprietary methodologies and resources.
Efforts to bolster LLM reasoning have explored diverse strategies. Process-based reward models (Uesato et al. 2022; Lightman et al. 2023a; Wang et al. 2023) guide models toward structured problem-solving, while RL approaches (Kumar et al. 2024) optimize performance through feedback-driven learning. Search algorithms, such as Monte Carlo Tree Search (MCTS) and Beam Search, have also been employed to enhance reasoning depth (Feng et al. 2024; Xin et al. 2024; Trinh et al. 2024). Although these methods have driven incremental gains, they fall short of the general reasoning prowess exhibited by the o1 series. Recently, the DeepSeek-R1 model (DeepSeek-AI 2025) has emerged as a competitive alternative, utilizing RL with the Group Relative Policy Optimization (GRPO) algorithm. Built on the 671-billion-parameter DeepSeek-V3, DeepSeek-R1 matches o1’s reasoning performance (DeepSeek-AI 2025). Yet, the sheer scale and computational demands of such models - often exceeding hundreds of billions of parameters - render them impractical for self-hosting by most organizations outside major technology firms, limiting their broader adoption.
In contrast, small LLMs, typically ranging from 1 to 10 billion parameters, present a resource-efficient alternative with potential for widespread deployment. Previous studies have demonstrated the feasibility of enhancing small LLMs through RL-based fine-tuning inspired by DeepSeek-R1 (Luo et al. 2025; Team 2025b). However, these efforts often rely on expansive datasets (hundreds of thousands to millions of samples) or incur significant computational costs, undermining their accessibility for resource-constrained settings. This tension motivates two central research questions:
@@
Our findings illuminate the promise of RL-based methods to enhance small LLMs’ reasoning capabilities, achieving competitive performance with minimal resources (Figure˜1). Simultaneously, they reveal critical challenges - such as data efficiency, optimization stability, and length constraints - that must be addressed to fully realize this potential. These insights lay the groundwork for developing lightweight, reasoning-capable LLMs suitable for resource-constrained environments, advancing the democratization of advanced AI technologies.
@@ 2 Methodology
High-Quality Dataset Curation
To minimize training costs while maximizing reasoning performance, we curate a compact, high-quality dataset tailored to mathematical reasoning. This dataset is derived from two existing sources: the s1 dataset (Muennighoff et al. 2025) and the DeepScaleR dataset (DeepSeek-AI 2025). By filtering and refining these datasets, we ensure that our training data is both relevant and challenging, enabling efficient learning for small LLMs.
@@
The s1 dataset (Muennighoff et al. 2025) is a general-purpose reasoning corpus comprising 59,029 questions sourced from diverse domains, including NuminaMATH (LI et al. 2024), AIME problems (1983–2021), OlympicArena (Huang et al. 2024), OmniMath (Gao et al. 2024), AGIEval (Zhong et al. 2023), probability questions from Stanford University’s Statistics Department PhD Qualifying Exams (https://statistics.stanford.edu), and brain-teasers from PuzzledQuant (https://www.puzzledquant.com). Although the dataset spans multiple disciplines - such as Astronomy, Biology, Chemistry, Computer Science, Geography, Mathematics, and Physics - our focus is exclusively on mathematical reasoning.
@@
The DeepScaleR dataset (Luo et al. 2025) contains 40,315 mathematics-specific questions drawn from AIME (1984–2023), AMC (prior to 2023), Omni-MATH, and the Still dataset. Unlike the s1 dataset, DeepScaleR is pre-filtered to focus solely on mathematics, with redundant questions removed and solutions extracted from raw text using retrieval-augmented generation (RAG) and advanced LLMs like Gemini-1.5-Pro-002. To further refine this dataset, we apply Qwen2.5-Math-7B-Instruct to exclude easy questions, reducing the set to 21,044 examples – open-deepscaler dataset. We opt for Qwen2.5-Math-7B-Instruct over DeepSeek-R1-Distill-Qwen-1.5B - used for the s1 dataset - to introduce diversity in filtering criteria and avoid excessive overlap between the two datasets.
@@
Reinforcement Learning Algorithm
To train small LLMs efficiently, we adopt the Group Relative Policy Optimization (GRPO) algorithm (Shao et al. 2024), as utilized in (DeepSeek-AI 2025). GRPO eliminates the need for a separate critic model - typically as large as the policy model - by estimating baselines from group scores, thereby reducing computational overhead. For each question $q$, GRPO samples a group of $G$ outputs $\{o_{1},o_{2},\dots,o_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$ and optimizes the policy $\pi_{\theta}$ by maximizing the following objective:
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{[q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)]}$
[…]
$\quad\quad\left.\left.\text{clip}\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)},1-\epsilon,1+\epsilon\right)A_{i}\right)\right)$
$\quad-\beta\mathbb{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})$ (1)
@@ 3 Experiments
To address the research questions outlined in Section 1 - namely, how reinforcement learning (RL) can enhance the reasoning abilities of small large language models (LLMs) and what practical insights emerge under computational constraints - we design three experiments to analyze the training behavior of small LLMs. These experiments aim to provide empirical evidence of performance improvements and offer actionable guidance for future research and industrial applications.
Experimental Setup
We select DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI 2025) as our base model for training. This 1.5-billion-parameter model, distilled from larger architectures, is chosen for its balance of efficiency and reasoning potential. Notably, we bypass the supervised fine-tuning (SFT) phase - typically a precursor to RL for performance enhancement (Chu et al. 2025) - hypothesizing that the model’s pretraining is sufficient to leverage RL directly. For the RL phase, we employ the Group Relative Policy Optimization (GRPO) algorithm, as detailed in Section 2, due to its computational efficiency.
Training is conducted on a cluster of 4 NVIDIA A40 GPUs (48GB VRAM each), imposing constraints that limit us to sampling 6 outputs per step with a maximum completion length of 4096 tokens. To facilitate this, we adapt open-r1 (Face 2025), an open-source reproduction of DeepSeek-R1 by the Hugging Face team, customizing it to align with our objectives. The training phase is restricted to 1 epoch, completed within a 24-hour window, reflecting real-world resource limitations. Hyperparameters and additional configurations are detailed in Appendix E.
Benchmark Datasets
To evaluate the reasoning capabilities of our small LLM, we choose five mathematics-focused benchmark datasets: AIME24  [^1] https://huggingface.co/datasets/AI-MO/aimo-validation-aime, MATH-500 (Lightman et al. 2023b; Hendrycks et al. 2021), AMC23  [^2] https://huggingface.co/datasets/AI-MO/aimo-validation-amc, Minerva (Lewkowycz et al. 2022a) and OlympiadBench (He et al. 2024). Details of the datasets are provided in Appendix˜C.
Baseline Models
To contextualize our results, we compare our trained model against a range of baselines: Llama-3.1-70B-Instruct (AI 2024a), o1-preview (AI 2024b), Qwen-2.5-Math-7B-Instruct (Yang et al. 2024), rStar-Math-7B (Guan et al. 2025), Eurus-2-7B-PRIME, Qwen2.5-7B-SimpleRL (Zeng et al. 2025) (Cui et al. 2025), DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI 2025), DeepScaleR-1.5B-Preview (Luo et al. 2025), Still-3-1.5B-Preview (Team 2025b).
This selection enables a robust comparison across model sizes, training methodologies, and reasoning strategies, highlighting the efficacy of our approach for small LLMs. Details of the baselines are provided in Appendix˜D.
Evaluation Metric
@@
Experiment 1: Impact of High-Quality Data.
In Experiment 1, we train the DeepSeek-R1-Distill-Qwen-1.5B model using the open-s1 dataset (18,615 samples) from Section 2, with a maximum completion length of 4096 tokens. We employ accuracy and format rewards, as described in Section 2. Although the full dataset corresponds to approximately 1500 global steps for one epoch, computational constraints (24-hour limit on 4x A40 GPUs) restrict training to 500 global steps.
@@
Experiment 2: Balancing Easy and Hard Problems.
Building on Experiment 1, we hypothesize that mixing easier problems with challenging ones could stabilize training and reduce completion lengths. We construct a dataset of 7000 samples: 3000 from open-s1, 3000 from open-deepscaler, and 1000 easier problems from the raw DeepScaleR dataset (Section 2). The maximum completion length is reduced to 3584 tokens, retaining accuracy and format rewards.
@@
Experiment 3: Controlling Length with Cosine Reward.
Experiment 3 uses the same 7000-sample dataset as Experiment 2, but replaces the accuracy reward with a cosine reward to better control output length, as outlined in Section 2. We also append an instruction to the system prompt: “Reply in English only, do not use other languages”, avoiding a computationally expensive language reward function. The maximum completion length remains 3584 tokens.
@@
Overall Comparison.
We select checkpoints at 100, 50, and 50 global steps from Experiments 1, 2, and 3, naming them Open-RS1, Open-RS2, and Open-RS3 (R for Reasoning, S for Small), respectively. These are evaluated against baselines from Section 3 across benchmarks from Section 3, using zero-shot pass@1 (Table 1).
@@ 1.5B Models
Table 1: Zero-shot pass@1 performance across benchmarks. Bold indicates the highest score per benchmark. Dashes (–) denote unavailable official scores. Scores for o1-preview are sourced from (AI 2024b); others from (Zeng et al. 2025; Luo et al. 2025). Our models are evaluated using the lighteval package (Fourrier et al. 2023).
@@
We further compare training costs [^3] Cost estimates are calculated based on pricing from https://www.runpod.io/pricing. and data efficiency (Tables 2 and 3, and Figure˜1). Our approach, using 7000 samples with 6 outputs per step (42,000 total samples), costs approximately \$42 on 4x A40 GPUs over 24 hours. In contrast, 7B models like Qwen2.5-7B-SimpleRL (\$1633) and Eurus-2-7B-PRIME (\$1088) and 1.5B models like DeepScaleR-1.5B-Preview (\$3629) and Still-3-1.5B-Preview (\$2268) require significantly more resources and data (e.g., 40k × 16 samples for DeepScaleR).